OpenStudy (anonymous):

Simplify the radical expression: √(18a2)

OpenStudy (anonymous):

Is that \[ \sqrt{18a^2} \] if it is you can rewrite it as \[\sqrt{9*2a^2} = 3a\sqrt{2}\]

OpenStudy (anonymous):

Thanks so much!! <(^_^)>